A marine biologist measures the body length of a bottlenose dolphin to be 343 cm. If the accepted value for the length of the said marine animal is 345 cm, what is the percentage error of the marine biologist/s measurement?


1
Expert's answer
2021-10-25T03:49:55-0400

345-343= 2

= 2/345× 100= 0.58%